# Advances and Open Challenges in Federated Learning with Foundation Models

### v2

@@ IV-C1 Weighted Average
To the best of our knowledge, currently, there is not FL aggregation technique specifically tailored for FedFM. The prevailing practice has gravitated towards adopting the basic vanilla versions of FedAvg or FedSGD, with the latter being equivalent to FedAvg when the number of local epochs ($E$) is set to one. Table V provides a summary of the aggregation methods adopted by existing FedFM works.
TABLE V: FL Aggregation Methods
Methods Aggregation
FedPAQ [30] FedSGD
SoteriaFL [38] FedSGD
FwdLLM [47] FedSGD
HeteroFL [31] FedAvg
LotteryFL [32] FedAvg
FjORD [33] FedAvg
H-FL [34] FedAvg
PruneFL [35] FedAvg
FedPM [36] FedAvg
FedTiny [37] FedAvg
FedPrompt [39] FedAvg
FedBERT [40] FedAvg
FedCLIP [41] FedAvg
FedPEFT [42] FedAvg
FedPETuning [43] FedAvg
SLoRA [44] FedAvg
FedOBD [45] FedAvg
FedIT [46] FedAvg
@@ IV-C2 Promising New Techniques
On the other hand, the potential for more sophisticated aggregation strategies remains. The emergence of Mixture of Experts (MoE)-based language models, exemplified by GLaM [57] and ST-MoE [58], has showcased exceptional capabilities across various tasks. These MoE architectures incorporate a myriad of smaller, specialized sub-models orchestrated by a routing function that aligns with the principle of dedicated local models in FL. The efficacy of MoE models is significantly influenced by their learned routing mechanisms [59]. This concept resembles the aggregation strategies in FL, where local model outputs are synthesized to form a global model. The routing mechanism in MoE could inspire a novel class of aggregation strategies in FL, where dynamic data-driven approaches could allocate model contributions in a similar manner. Exploring aggregation strategies that mirror MoE’s routing scheme, adapting it to the distributed nature of FL, could inspire new FedFM model aggregation methods. Such exploration might involve designing algorithms that not only learn from local data but also intelligently decide how to combine local updates to form a robust global model, potentially redefining FedFM aggregation approach as a learned component rather than a fixed rule.
@@ IV-D Computationally Efficient FedFM
This section delves into recent advances in training FMs via FL techniques, with a focus on enhancing computational efficiency. Specifically, we examine a set of methods including: 1) Parameter-Efficient Fine-Tuning (PEFT), 2) Prompt Tuning, and 3) Instruction Tuning. These three types of methods have been shown to be feasible.
@@ IV-D1 Parameter-Efficient Fine-Tuning
PEFT refines FMs by optimizing a minimal subset of model parameters, strategically positioned throughout the model. This approach significantly reduces computational and storage requirements, demonstrating that even large models can be efficiently adapted by tweaking a handful of parameters [60]. A pioneering PEFT technique, adapter tuning [61], modifies pre-trained models with minimal parameter adjustments. It integrates specialized adapters with a bottleneck design between model layers. BitFit [62] selectively updates only the bias terms within FMs, leaving other components unchanged. Low Rank Adaptation (LoRA) [63] updates attention weights through low-rank matrices, thereby minimizing the number of parameters required for training. Typically, the proportion of trainable parameters in such methods is around 1% of the entire FM. For instance, conventional fine-tuning of a model like GPT-3 necessitates adjusting approximately 175 billion parameters—a task that is untenable for most industry and academic settings. By employing LoRA and focusing solely on low-rank matrices within each transformer layer, the training process involves just 37.7 million parameters.
@@
Under FL settings, Sun et al. [42] introduced FedPEFT, an innovative approach for fine-tuning FMs. This method freezes the majority of the model weights, focusing on adjusting a minimal subset of parameters tailored to specific downstream tasks. FedPEFT evaluates the efficiency of four distinct PEFT strategies including, Head-tuning, Bias, Adapter and Prompt-in the context of the Vision Transformers (ViT-B) with 85 million parameters. Their findings demonstrate that these targeted fine-tuning methods achieve comparable results to centralized model fine-tuning with non-IID data, while remarkably reducing communication overhead by over 99%. FedCLIP [12] also experimented with an Adaptor-based PEFT method using adapter aggregation even based on FedAvg, showing the better performance than standard FedProx and FedAvg. FedPETuning [43] conducted an extensive benchmark analysis of adapter-based fine-tuning techniques under FL settings. Their findings indicate that both approaches, integrating additional adapters and utilizing LoRA, can achieve comparable accuracy. However, LoRA-based fine-tuning incurs only one third of the communication overhead required by methods involving extra adapters.
The above methods are primarily designed for IID scenarios. Compared to complete fine-tuning, they often result in lower accuracy. In contrast, SLoRA [44] introduced an optimization of LoRA tailored for non-IID FL settings. It employs a ratio to moderate the influence of updates from individual clients, aiming to counteract the drift caused by heterogeneous data distributions. Although it reportedly matches the performance levels of full model fine-tuning, it requires a considerable warm-up period to ensure that all clients share a common initialization of LoRA.
@@
FwdLLM [47] is designed to adapt FMs for use on mobile devices with stricter resource constraints. To address this issue, it integrates Backpropagation (BP)-free training with parameter-efficient training approaches. Rather than relying on traditional back-propagation to calculate precise gradients, the BP-free method introduces minor and self-generating perturbations to these ML model parameters. It then evaluates how these perturbations affect model predictions compared to the original unperturbed model. If a perturbation results in predictions that are closer to the actual labels, it is considered to be guiding the model towards the global optimum. Remarkably, FwdLLM demonstrates substantial performance enhancements, achieving convergence three orders of magnitude faster and reducing the memory footprint by close to 15 times.
@@ IV-E Communication Efficient FedFM
The complexity of FMs directly translates to inflated communication overhead between FL clients and the FL server in FedFM, hindering efficient collaboration. Although as discussed in Section IV-D, it is not always necessary to transmit an entire FM, fine-tuning billion-parameter FMs by adapter still requires the transmission of millions of parameters. This highlights the need for communication efficient FedFM research. In this section, we highlight two primary strategies for enhancing efficiency: 1) model pruning and 2) model compression.
@@ V Towards Trustworthy FedFM
Attacks and defenses on traditional FL process have been extensively studied [70]. In this section, we review selected studies on robustness and privacy issues in FL, and point out the potential new changes or challenges FedFM could bring about. We further break down each domain into specific areas as shown in Fig. 3.
Figure 3: Taxonomy for trustworthy FedFM techniques consisting of two main domains: 1) Robustness and 2) Privacy, each of which is further divided into specific strategies that address the challenges of FedFM.
@@ V-A1 Poisoning Attacks
FL poisoning attacks aim to compromise either the global model or the training process. They can be divided into two categories: 1) untargeted attacks and 2) targeted attacks, based on whether the attacks aim to disrupt global model convergence or manipulate the model outputs.
@@
However, achieving the objective of such attacks can be challenging in FedFM, due to the large sizes and high local heterogeneity of FM training tasks. Firstly, to achieve validation loss maximization, an attacker needs to repeatedly train a local shadow model to form a differentiable poisoning sample optimization objective. This is infeasible in FedFM considering the huge costs introduced by FMs. Secondly, the local ML tasks can be heterogeneous, making it hard for such attacks to compromise global FM convergence or performance.
@@
Defending against targeted attacks is more challenging under FedFM settings. Specifically, due to the high complexity of FM training tasks, the poisoned input that triggers the FM misbehavior is more stealthy. Existing backdoor detection approaches [84, 85] mainly rely on generating cross-category sample transfer shortcuts via optimization, given the model and total number of categories. However, the ML tasks for FMs introduce a enormously high level of complexity (e.g., millions of image classification categories), making existing defense mechanism against targeted FL attacks intractable.
V-A2 Byzantine-Robust Local Model Aggregation
@@
Other Hybrid Schemes: Research works have emerged [91, 92, 93] in an attempt to form hybrids schemes to take advantages of the aforementioned by combining them together. These works often update a benchmark model with a central clean dataset, and lower the weights assigned to local model updates that significantly deviate from the benchmark (or even discard them completely) during aggregation. Although these methods have achieved superior robustness, they have also inherited the limitations from the aforementioned categories of approaches which make them incompatible with the FedFM settings.
@@ V-B Privacy of FedFM
Privacy preservation has always been an important focus of FL research. A wide range of privacy attacks have been studied and potential defenses have been proposed.
@@ V-B1 Privacy Attacks
Privacy attacks in FL are generally designed to access the victim client’s private information given based on auxiliary information essential for FL model training (e.g., the victim’s local model updates or gradients). They can be divided into two categories: 1) membership inference attacks, which attempt to infer the involvement of a specific data sample in FL training, and 2) data reconstruction attacks, which attempt to recover a victim client’s original training data.
@@ VI Towards Incentive Mechanisms for FedFM
The remarkable success of FL depends on clients (a.k.a., data owners) actively engaging during the training process. In reality, data owners might hesitate to join FL without proper compensation, especially if it involves significant commitment of local resources for training FMs. Thus, developing robust incentive mechanisms is imperative to motivating clients to participate in the FedFM training process, while deterring misbehaviours through punitive measures. Table VI provides an overview of existing incentive mechanisms in FL.
@@ VI-A FL Incentive Mechanisms
TABLE VI: Summary of the Main Categories of Incentive Mechanisms in FL
@@ VI-B Discussions on FedFM Incentive Mechanisms
Figure 4: Illustration of challenges faced by incentive mechanism design in FedFM.
@@ VI-C FedFM Participant Selection
The model selection process in FedFM encompasses two critical dimensions: 1) horizontal fitness for task-specific training and 2) vertical capability for training.
@@ VI-C2 Vertical Capability for Training
In [158, 159], FedLEGO was proposed to address model heterogeneity issues in conventional FL. It treats participants’ local models as a LEGO toy, disassembling them into bricks by layers. Subsequently, FedLEGO reassembles these bricks into new model structures. Drawing inspiration from heterogeneous FL, we envision that FM construction can be performed similar to LEGO, where it can consist of various functional building blocks hosted by diverse FedFM participants to be assembled together. We envision a multi-step approach for FedFM model training through layer-wise decomposition, function-wise layer grouping, reassembly candidate generation, and candidate stitching.
  • Layer-wise Decomposition: Initially, the FedFM model is disassembled into layers, focusing on the operation type of each model component. This step aims to identify the layers and their corresponding operation types within the FMs.

### v4

@@ IV-C1 Weighted Average
Figure 5: Illustration of weight averaging strategy of aggregation methods in FedFM.
To the best of our knowledge, currently, there is no FL aggregation technique specifically tailored for FedFM. The prevailing practice has gravitated towards adopting the basic vanilla versions of FedAvg or FedSGD, with the latter being equivalent to FedAvg when the number of local epochs ($E$) is set to one, as shown in Fig. 5. Table III provides a summary of the aggregation methods adopted by existing FedFM works.
@@ IV-C2 Mixture of Experts
On the other hand, the potential for more sophisticated aggregation strategies remains. The emergence of large language models based on Mixture of Experts (MoE)[75], exemplified by GLaM [76] and ST-MoE [77], has demonstrated exceptional capabilities across various tasks. These MoE architectures incorporate a myriad of smaller, specialized sub-models orchestrated by a routing function (gating network) that aligns with the principle of dedicated local models in FL, as shown in Fig. 6. The efficacy of MoE models is significantly influenced by their learned routing mechanisms [78]. This concept resembles the aggregation strategies in FL, where local model outputs are synthesized to form a global model. The routing mechanism in MoE could inspire a novel class of aggregation strategies in FL, where dynamic data-driven approaches could allocate model contributions in a similar manner. Exploring aggregation strategies that mirror MoE’s routing scheme and adapting it to the distributed nature of FL, could inspire new FedFM model aggregation methods. Such exploration might involve designing algorithms that not only learn from local data, but also intelligently decide how to combine local updates to form a robust global model, potentially redefining the FedFM aggregation approach as a learned component rather than a fixed rule.
Figure 6: MoE strategies of FedFM aggregation methods.
MoE-FL [79] uses MoE to aggregate a robust global model on the server. It assumes trust between the server and clients where each client shares partial data with the server. Although the produced robust aggregation model can filter outliers (poisoned/corrupted/outdated client models), sharing partial local data may be prohibited in most FL applications with privacy requirements. To address the data heterogeneity problem in FL, FedMix [80] allows each client to utilize MoE to adaptive select which client expert models are relevant to it. Fed-MoEs [81] obtain all client local models pre-trained on local data. These pre-trained models, the aggregated global model, and the private client model form the experts of MoE on each client. The client uses MoE to dynamically combine these experts for domain adaptation, addressing the domain shift problem. However, it introduces significant communication overhead due to the need to transmit multiple models to each client.
PFL-MoE [82] combines MoE with personalized FL (PFL) to improve model personalization, while maintaining model generalization. It constructs an MoE for each client, involving the gating network and two expert models - the global model and the local personalized model. Each client uses the gating network to generate two weights for the outputs of the two expert models, and mixes the weighted outputs as the final output. pFedMoE [83] first introduced MoE into model-heterogeneous personalized FL (MHPFL) to enable collaboration among heterogeneous client models with efficient communication and computation. Each client is constructed with an MoE formed by a gating network, the shared homogeneous feature extractor as the global expert, and the heterogeneous client model as the local expert. For each local data sample, the gating network produces a pair of personalized weights to represent the two experts. The client performs weighted averaging on them to form a complete representation containing both generalized and personalized features.
@@ IV-D Computationally Efficient FedFM
This section delves into recent advances in training FMs via FL techniques, with a focus on enhancing computational efficiency. Specifically, we examine a set of methods including: 1) Parameter-Efficient Fine-Tuning (PEFT), 2) Prompt Tuning, and 3) Instruction Tuning. These three types of methods have been shown to be feasible, as shown in Fig. 7.
Figure 7: Methods for enhancing FedFM computational efficiency.
@@ IV-D1 Parameter-Efficient Fine-Tuning
PEFT refines FMs by optimizing a minimal subset of model parameters, strategically positioned throughout the model. This approach significantly reduces computational and storage requirements, demonstrating that even large models can be efficiently adapted by tweaking a handful of parameters [84].
@@
A pioneering PEFT technique, adapter tuning [85], modifies pre-trained models with minimal parameter adjustments. It integrates specialized adapters with a bottleneck design between model layers. BitFit [86] selectively updates only the bias terms within FMs, leaving other components unchanged. Low-Rank Adaptation (LoRA) [87] updates attention weights through low-rank matrices, thereby minimizing the number of parameters required for training. Typically, the proportion of trainable parameters in such methods is around 1% of the entire FM. For instance, conventional fine-tuning of a model like GPT-3 necessitates adjusting approximately 175 billion parameters—a task that is untenable for most industry and academic settings. By employing LoRA and focusing solely on low-rank matrices within each transformer layer, the training process involves just 37.7 million parameters.
@@
Under FL settings, Sun et al. [48] introduced FedPEFT, an innovative approach for fine-tuning FMs. This method freezes the majority of the model weights, focusing on adjusting a minimal subset of parameters tailored to specific downstream tasks. FedPEFT evaluates the efficiency of four distinct PEFT strategies including, Head-tuning, Bias, Adapter and Prompt-in the context of the Vision Transformers (ViT-B) with 85 million parameters. Their findings demonstrate that these targeted fine-tuning methods achieve comparable results to centralized model fine-tuning with non-IID data, while remarkably reducing communication overhead by over 99%. FedCLIP [12] also experimented with an Adaptor-based PEFT method using adapter aggregation even based on FedAvg, showing the better performance than standard FedProx and FedAvg.
FedPETuning [49] conducted an extensive benchmark analysis of adapter-based fine-tuning techniques under FL settings. Their findings indicate that both approaches, integrating additional adapters and utilizing LoRA, can achieve comparable accuracy. However, LoRA-based fine-tuning incurs only one-third of the communication overhead required by methods involving extra adapters. SLoRA [50] introduced an optimization of LoRA tailored for non-IID FL settings. It employs a ratio to moderate the influence of updates from individual clients, aiming to counteract the drift caused by heterogeneous data distributions. Although it reportedly matches the performance of full model fine-tuning, it requires a considerable warm-up period to ensure that all clients share a common initialization of LoRA. pFedLoRA [88] first introduced LoRA into model-heterogeneous FL. It consists of a lightweight linear low-rank adapter shared by clients with heterogeneous local models for effective knowledge exchange, achieving efficient communication and computation. SA-FedLoRA [89] applied LoRA to fine-tune a large-scale pre-trained model in FL. It freezes the pre-trained model and trains a LoRA adapter with reduced ranks as communication rounds increase, improving model performance while incurring low communication and computational overheads. FFA-LoRA [90] utilizes LoRA to fine-tune LLMs. It freezes the pre-trained LLMs and the randomly initializes the non-zero matrix of LoRA. It only fine-tunes the zero-initialized matrix of LoRA to address data heterogeneity, enhance privacy preservation under differential privacy (DP), and improve computational efficiency.
Cross-device FL systems often include devices with diverse resource capacities, leading to differences in model training efficiency [91]. To mitigate this issue, several methods have been proposed to tailor model architectures for resource-heterogeneous FL systems. LoRA-based FedPEFT, for example, offers notable flexibility and adaptability in fine-tuning frozen FMs without overloading client devices. [92] suggested assigning LoRA adapters to different numbers of layers for clients based on a randomly generated mask matrix. Alternatively, selecting diverse LoRA ranks according to the system capabilities of clients is another strategy. FlexLoRA [93] dynamically adjusts local LoRA ranks. It reconstructs the full-sized LoRA module for server-side aggregation, followed by SVD-based parameter redistribution. However, [94] observed that this reconstruction-redistribution approach results in performance degradation compared to homogeneous LoRA. Consequently, they proposed HETLORA [94] which employs zero-padding to standardize module sizes before aggregation, and then truncates the global LoRA modules to fit the specific rank requirements of subsequent clients.
@@
FwdLLM [53] is designed to adapt FMs for use on mobile devices with stricter resource constraints. To address this issue, it integrates Backpropagation (BP)-free training with parameter-efficient training approaches. Rather than relying on traditional BP to calculate precise gradients, the BP-free method introduces minor and self-generating perturbations to these ML model parameters. It then evaluates how these perturbations affect model predictions compared to the original unperturbed model. If a perturbation results in predictions that are closer to the actual labels, it is considered to be guiding the model towards the global optimum. Remarkably, FwdLLM demonstrates substantial performance enhancements, achieving convergence three orders of magnitude faster and reducing the memory footprint by close to 15 times.
@@ IV-E Communication Efficient FedFM
The complexity of FMs directly translates to inflated communication overhead between FL clients and the FL server in FedFM, hindering efficient collaboration. Although as discussed in Section IV-D, it is not always necessary to transmit an entire FM, fine-tuning billion-parameter FMs by adapter still requires the transmission of millions of parameters. This highlights the need for communication efficient FedFM research. In this section, we highlight two primary strategies for enhancing efficiency: 1) model pruning and 2) model compression, the readers can refer to Fig. 4.
@@ V Towards Trustworthy FedFM
Figure 8: Trustworthy FedFM research works can be divided into three domains: 1) Robustness, 2) Privacy, and 3) Intellectual Property Protection, each of which is further divided into specific strategies that address the challenges of FedFM.
Attacks and defenses on traditional FL processes have been extensively studied [101]. In this section, we review selected studies on robustness and privacy issues in FL, and point out the need for intellectual property protection in FedFM. We further break down each domain into specific areas as shown in Fig. 8.
@@ V-A1 Poisoning Attacks
FL poisoning attacks aim to compromise either the global model or the training process. They can be divided into two categories: 1) targeted attacks and 2) untargeted attacks, based on whether the attacks aim to manipulate model outputs or disrupt global model convergence (Fig. 9).
@@
V-A2 Poisoning Defenses
Defense for Targeted Attacks: Defending against targeted attacks is more challenging under FedFM settings. Specifically, due to the high complexity of FM training tasks, the poisoned input that triggers the FM misbehavior is more stealthy. Existing backdoor detection approaches [115, 116] mainly rely on generating cross-category sample transfer shortcuts via optimization, given the model and total number of categories. However, the ML tasks for FMs introduce a enormously high level of complexity (e.g., millions of image classification categories), making existing defense mechanism against targeted FL attacks intractable.
Defense for Untargeted Attacks: However, achieving the objective of such attacks can be challenging in FedFM, due to the large sizes and high local heterogeneity of FM training tasks. Firstly, to achieve validation loss maximization, an attacker needs to repeatedly train a local shadow model to form a differentiable poisoning sample optimization objective. This is infeasible in FedFM considering the huge costs introduced by FMs. Secondly, the local ML tasks can be heterogeneous, making it hard for such attacks to compromise global FM convergence or performance.
@@ V-A2 Poisoning Defenses
Hybrid Schemes: Research works have emerged [122, 123, 124] in an attempt to form hybrids schemes to take advantages of the aforementioned by combining them together. These works often update a benchmark model with a central clean dataset, and lower the weights assigned to local model updates that significantly deviate from the benchmark (or even discard them completely) during aggregation. Although these methods have achieved superior robustness, they have also inherited the limitations from the aforementioned categories of approaches which make them incompatible with the FedFM settings.
Figure 9: Illustration of poisoning and privacy attacks on FedFM.
@@ V-B Privacy of FedFM
Privacy preservation has always been a central focus of FL research. A wide range of privacy attacks have been studied and potential defenses have been proposed.
@@ V-B1 Privacy Attacks
Privacy attacks in FL are generally designed to access the victim client’s private information given based on auxiliary information essential for FL model training (e.g., the victim’s local model updates or gradients). They can be divided into two categories: 1) membership inference attacks, which attempt to infer the involvement of a specific data sample in FL training, and 2) data reconstruction attacks, which attempt to recover a victim client’s original training data (Fig. 9).
@@ V-B2 Privacy Defenses
Figure 10: Illustration of intellectual property protection strategies for FedFM.
V-C FedFM Intellectual Property Protection
Intellectual property protection is crucial for safeguarding FedFM against unauthorized use (e.g., model theft [146]). In the context of FedFM, we explore two main intellectual property protection strategies: 1) black-box fine-tuning, and 2) watermarking (Fig. 10). By integrating these intellectual property protection strategies, FedFM can ensure the secure and efficient use of FMs across distributed networks, making FedFM more robust and secure for collaborative AI model development
Black-Box Fine-Tuning: Black-box fine-tuning refers to methods that fine-tune FMs without accessing the model’s internal parameters [147]. This approach is particularly relevant in FL, where maintaining the integrity and ownership of the original models is crucial. Black-box fine-tuning methods often add new parameters while keeping the original model unchanged, thus preserving the original model’s intellectual property. For example, Fed-BBPT [148] is a prompt tuning framework that enables the joint training of a lightweight prompt generator across multiple clients. This framework allows clients to fine-tune their models locally without accessing or altering the core model parameters. Similarly, the method in [149], called FedBPT, uses an evolutionary algorithm to train optimal prompts that enhance the performance of frozen FMs. These black-box fine-tuning methods ensure that local fine-tuning of FMs can be achieved with robust intellectual property protection. However, current research mainly focuses on few-shot learning with small datasets for FM fine-tuning, indicating a need for further exploration with larger datasets and multiple data types to fully leverage the potential of black-box fine-tuning for FedFM.
Watermarking: Watermarking is a widely-adopted technique for intellectual property protection by embedding identifiers into models to prove ownership. In FL environments, [146] introduced WAFFLE, a solution for embedding watermarks into global models. This technique ensures that the ownership of the model can be verified even when the model is distributed across multiple clients. More recently, [150] developed DUW, which embeds a unique key into each client’s local model. This key can identify the source of a leaked model and verify ownership, thereby providing a robust mechanism for intellectual property protection in federated settings. The use of watermarking in FedFM not only can secure the global model, but also can protect individual client models, thereby ensuring comprehensive intellectual property protection across the FedFM network.
@@ VI Towards Incentive Mechanisms for FedFM
The remarkable success of FL depends on clients (a.k.a., data owners) actively engaging during the training process. In reality, data owners might hesitate to join FL without proper compensation, especially if it involves significant commitment of local resources for training FMs. Thus, developing robust incentive mechanisms is imperative to motivating clients to participate in the FedFM training process, while deterring misbehaviours through punitive measures. Table V provides an overview of existing incentive mechanisms in FL.
@@ VI-A FL Incentive Mechanisms
TABLE V: Summary of the Main Categories of Incentive Mechanisms in FL
@@ VI-A2 Game Theory-based Methods
Figure 11: FedFM incentive mechanisms can be divided into three domains: 1) Participant Selection, 2) Reward Distribution, and 3) Contribution Evaluation, each of which is further divided into specific strategies that address the challenges of FedFM.
@@ VI-C FedFM Participant Selection
Figure 12: Illustration of participant selection in FedFM.
Fig. 12 summarizes FedFM participants selection. The model selection process in FedFM encompasses two critical dimensions: 1) horizontal fitness for task-specific training and 2) vertical capability for training.
@@ VI-C2 Vertical Capability for Training
In [194, 195], FedLEGO was proposed to address model heterogeneity issues in conventional FL. It treats participants’ local models as a LEGO toy, disassembling them into bricks by layers. Subsequently, FedLEGO reassembles these bricks into new model structures. Drawing inspiration from heterogeneous FL, we envision that FM construction can be performed similar to LEGO, where it can consist of various functional building blocks hosted by diverse FedFM participants to be assembled together. We envision a multi-step approach for FedFM training through layer-wise decomposition, function-wise layer grouping, reassembly candidate generation, and candidate stitching.
  • Layer-wise Decomposition: Initially, the FedFM is disassembled into layers, focusing on the operation type of each model component. This step aims to identify the layers and their corresponding operation types within the FMs.